kittybasil:

After he made deposits of $85 and $75, a student's account was still $55 overdrawn. What was his checking account balance (in dollars) before the deposits?

kittybasil:

Just checking in case I read the language wrong. Here's what I did:\[-55-85-75=-215\]as in: $215 dollars overdrawn

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kittybasil Just checking in case I read the language wrong. Here's what I did:\[-55-85-75=-215\]as in: $215 dollars overdrawn \(\color{#0cbb34}{\text{End of Quote}}\) correct - good job - congrats !

jhonyy9:

@supie opinion pls. ? ty

AZ:

That's correct. You can think of the original overdrawn balance as 'x' Even after depositing 85 and 75 dollars, his account was still overdrawn by 55 dollars (so he had a balance of -55) So we have x + 85 + 75 = -55 And to solve for x, it would then be just like how you set it up x = -55 - 85 - 75 = -215 So good job!